The reaction between solid white phosphorus and oxygen produces solid tetraphosphorusdecoxide (P4O10)



A. Write the balanced chemical equation of the reaction


b. Determine the mass of tetraphosphorusdecoxide (P4O10) formed if 26.0 g of phosphorus (P4) and 49.5 g oxygen (O2) are combined.


C. How much of the excess reactant remains after the reaction stops.

1
Expert's answer
2022-04-27T09:41:01-0400

P4 + 5O2 → P4O10

n(P4)=n(P4O10)= 0.21 mol, the oxygen is in excess (is 0.1), m(P4O10)=0.21*155.90=32.739 g